OpenStudy (anonymous):

how do you find out if a fraction is proportional i know proportional are numbers that can be divided evenly such as 30 , 50 but i dont know with fractions

hero (hero):

There's basically two ways to do it. Piggybacking from what @raffle_snaffle posted, suppose you had two fractions: \[\frac{4}{16} = \frac{2}{8}\]

hero (hero):

Then you can either cross multiply to get: 32 = 32

hero (hero):

Or you can reduce both fractions and observe that they both reduce to 1/4

OpenStudy (anonymous):

i understand but what if it is a mixed number like 6 1/4

hero (hero):

In order to find out if two fractions are proportional to each other, you need TWO fractions.

hero (hero):

6 + 1/4 is only one fraction

OpenStudy (anonymous):

ok other ones 8 1/2

hero (hero):

So you basically do the same thing 6 + 1/4 = 8 + 1/2 1/4 - 1/2 = 8 - 6 1/4 - 2/4 = 2 -1/4 = 2 -1 = 8 False

hero (hero):

If you have any questions about how I did that, let me know

hero (hero):

Two fractions that are proportional to each other will always be equal. If they are not equal, then they are not proportional.

OpenStudy (anonymous):

ok thanks il observe this a bit

hero (hero):

Just let me know if you have any questions. I expect you to have a few.

OpenStudy (anonymous):

ok il try one on my own and see how i do here

hero (hero):

Can you post it here please so that I know what you are working on?

hero (hero):

You can still do it on your own and we can discuss it afterward

OpenStudy (anonymous):

6 1/4 and 7 1/5 1/4 - 1/5=1/20 7 - 6 =1 so 1/20 and 1 are not proportional so no however that last part u did its out of nowhere so i didnt do it (-1 =8)

OpenStudy (anonymous):

ok i think u divided

hero (hero):

Make sure you do it this way: First set the fractions equal to each other. That's the first step

hero (hero):

When you finish, you need to have an integer on one side and an integer on the other side.

hero (hero):

The last step is to cross multiply

hero (hero):

6 1/4 = 7 1/5 1/5 - 1/4 = 7 - 6 1/20 = 1 Cross multiply 1 = 20 False

hero (hero):

Do you get it now?

OpenStudy (anonymous):

yep thank u

hero (hero):

Hang on a minute. Let me find one for you to do

hero (hero):

Nevermind, I think you get the hang of it. But chances are if the whole numbers are different, then they will not be equal. For example 3 + 4/7 = 3 + 8/14 That is proportional because 3 - 3 = 8/14 - 4/7 0 = 8/14 - 8/14 0 = 0

hero (hero):

But 4 + 1/4 = 5 + 3/4 That is not proportional

OpenStudy (anonymous):

hah ok thanks for the help